OpenStudy (anonymous):

I will fan and medal: 6(t - 2) = 2t + 2

OpenStudy (anonymous):

Simplifying 6(t + -2) = 2t + 2 Reorder the terms: 6(-2 + t) = 2t + 2 (-2 * 6 + t * 6) = 2t + 2 (-12 + 6t) = 2t + 2 Reorder the terms: -12 + 6t = 2 + 2t Solving -12 + 6t = 2 + 2t Solving for variable 't'. Move all terms containing t to the left, all other terms to the right. Add '-2t' to each side of the equation. -12 + 6t + -2t = 2 + 2t + -2t Combine like terms: 6t + -2t = 4t -12 + 4t = 2 + 2t + -2t Combine like terms: 2t + -2t = 0 -12 + 4t = 2 + 0 -12 + 4t = 2 Add '12' to each side of the equation. -12 + 12 + 4t = 2 + 12 Combine like terms: -12 + 12 = 0 0 + 4t = 2 + 12 4t = 2 + 12 Combine like terms: 2 + 12 = 14 4t = 14 Divide each side by '4'. t = 3.5 Simplifying t = 3.5